How does cos^2A+cos^2B+cos^2C+2cosAcosBcosC=1,where A,B,C are angles of triangles?

Grade:12

2 Answers

Rathod Shankar AskiitiansExpert-IITB
13 years ago

   Dear Aparna Tiwari,       

 

          2cosAcosBcosC =(2cosAcosB)cosC                               here A+B+C=pi=∏       thereforen  C=pi-(A+B)
                                =(cos(A + B) + cos(A - B))cosC
                                =cos(pi - C)cosC + cos(A - B)cos(pi - (A + B))
                                =-2cos^2C - 2cos(A + B)cos(A - B)

                                =-cos^2C - cos^2A + sin^2B

 

cos^2A+cos^2B+cos^2C+2cosAcosBcosC=cos^2A+cos^2B+cos^2C-cos^2C - cos^2A + sin^2B

                                                                              =   cos^2B+ sin^2B =1

 

 


Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar respectively : Click here to download the toolbar..



Askiitians Expert

Rathod Shankar Singh

IIT Bombay

 

Lab Bhattacharjee
6 years ago
As cos(A+B)=cos(pi-C)=-cos Cπ2
 
cos^2A+cos^2B+cos^2C=1+\cos^2A-\sin^2B+\cos^2C=1+cos(A+B)cos(A-B)+cos^2C=1-cos C cos(A-B)-cos C cos(A+B)
 
=1-cos C[cos(A+B)+cos(A-B)]=1-cos C(2ocs Acos B)